Columbia University Information Technology

Service Alerts

RESOLVED SERVICE ALERTS

Resolved Service Alerts

Resolved Service Alerts

  • Network - Ethernet
    (Approximate) 10/21/2008 - 02:35pm to 10/21/2008 - 03:30pm

    Today between 2:35pm and 3:30pm, technicians will be performing emergency maintenance in the network closet that services the 5th, 6th and 7th floors of 542 W 112th St.  This work is necessary in order to resolve an electrical issue at the site.

    Total downtime should be approximately 25 minutes.

     

  • Network - Ethernet and Wireless
    10/21/2008 - 08:00am to 10/21/2008 - 12:00pm

    CUIT is performing necessary maintenance on this service.   This scheduled maintenance is expected to be completed on-time.

    During the scheduled time of this Service Alert, The Chrystie Field House and the Wien Pressbox at Baker Field Athletics Complex will experience:

    -Loss of network connectivity

    CUIT apologizes for any inconvenience this may cause, but this action is necessary to ensure a more reliable service.  If clients continue to experience issues after the service-interruption hours, please contact the CUIT Helpdesk at:

    Web:     http://www.columbia.edu/cuit/support/expanded.html

    Phone:   212-854-1919

    Email:     askcuit@columbia.edu

     

    For a list of active and future Service Alerts, please visit http://www.columbia.edu/acis/alerts-list.html.

     

  • Central Server
    10/21/2008 - 03:00am to 10/21/2008 - 05:00am

    We will be performing emergency maintenance on the Central Groups servers tomorrow morning between 3:00 AM and 5:00 AM.  During this time shares on C-GP1, C-GP2, and C-GP3 and will be unavailable.  This maintenance is being done to add more disk space to the servers.

     

    Update 10/21/2008 at 7:05am:

    This work was completed successfully.

  • Central Server
    (Approximate) 10/20/2008 - 04:10pm to 10/20/2008 - 04:25pm

    The Central Groups servers are having a problem which is preventing people from being able to access them. These are the C-GP1 and C-GP2 network drives.  We are working on them now.  There is no ETA at this time.

     

    Update 10/20/2008 at 4:25pm:

    Service has been restored.  We had a cluster node hang and failover didn't happen automatically.  We are looking into it.

  • ReportWeb
    10/20/2008 - 12:13pm to 10/20/2008 - 01:29pm

    Users are experiencing problems while signing on to the Report.Web server. We are investigating and looking into it. There is no ETA as of this time.

     

    Update 10/20/2008 at 1:29pm:

    The problem was identified as database index corruption. The index was rebuilt and the application is now back online.